# BayeSN and SALT: A Comparison of Dust Inference Across SN Ia Light-curve Models with DES5YR

### v2

@@ 4.2 Impact of non-Ia Contamination
Here, we aim to test the impact of non-Ia contamination on dust inference using BayeSN. To do so, we simulate the Double $R_{V}$, Double $E(B-V)$ test with the addition of templates of non-Ia SNe from Vincenzi et al. (2021). We include four classes of non-Ia, following Vincenzi et al. (2024): SN Iax and SN Ia 91bg from Kessler et al. (2019) following revisions from Vincenzi et al. (2019), and core collapse SED templates from Vincenzi et al. (2019) with rates from Strolger et al. (2015) and Shivvers et al. (2017) with the modified Li et al. (2011) luminosity functions from Vincenzi et al. (2021). A full description of the non-Ia templates is available in Vincenzi et al. (2021). We classify our simulated SN light-curves using the SuperNNova classifier from Möller & de Boissière (2020), taking the trained SuperNNova model directly from the DES 5-year analysis, where a further review is available.
@@ 5.2 BayeSN Dust Estimates
The true simulated values input into the Dust2Dust simulations and our BayeSN posterior metrics across all 50 simulacra of our simulations are detailed in Tables 3 through 6.
@@
Firstly, considering $\mu_{R_{V}}$ across all simulated variations the BayeSN estimates are in very close agreement with the truth, within $1\sigma$ in many cases. The most notable discrepancy is $\bar{\mu}_{R}=2.12(0.02)$ for the $\mu_{R_{V}}=2.0$ high-mass bin in the case of the double $R_{V}$, $E(B-V)$ simulation with the Gaussian $\beta_{\rm SN}$ distribution, shown in Table 6. The Gaussian $\beta_{\rm SN}$ distribution has no equivalent in BayeSN, and was chosen to test potentially diverging choices in how the two models treat dust; our expectation was that this would be the simulated variant where BayeSN would perform the worst. A discrepancy of ${\sim}0.1$ in $\mu_{R_{V}}$ is small compared to the statistical uncertainty on real inferred $\mu_{R_{V}}$ values; this result could not explain previous disagreements between the findings of BayeSN and Dust2Dust. Additionally, the estimate of $\bar{\mu}_{R}$ in the low-mass bin for this simulation exactly matches the truth of 3.0, showing that this specific model mismatch does not cause particular issues for BayeSN. Our posteriors on $\mu_{R_{V}}$ are relatively well-calibrated considering the inherent model-mismatch in this analysis; $N_{95}$ is close to 95 per cent across all of our simulation variants for $\mu_{R_{V}}$.
@@
While $\sigma_{R}$ is recovered perfectly for our Double $R_{V}$, single $E(B-V)$ simulation, it is noticeable that BayeSN estimates of $\sigma_{R}$ are generally further than the true simulated values than for either $\mu_{R_{V}}$ or $\tau_{E}$. This difference is not large and could not explain historic differences between results of SALT-based analyses and BayeSN. For simulations with $\mu_{R_{V}}=3.0$, $\bar{\sigma}_{R}$ is typically overestimated, the most notable example being 0.77(0.04) for simulated $\sigma_{R}=0.5$. For simulations with $\mu_{R_{V}}=2.0$, however, $\bar{\sigma}_{R}$ is typically slightly underestimated, the most notable example being 0.41(0.02) for simulated $\sigma_{R}=0.5$. There are a few reasons why this might be the case. The dust contribution to the $c_{\rm int}$ term in the Dust2Dust simulations (as discussed in Section 2.3) may be impacting the inferred value of $\sigma_{R}$. Alternatively, this might result from the influence of the prior on $\sigma_{R}$ posteriors. $\sigma_{R}$ is typically less constrained by data than either $\mu_{R_{V}}$ or $\tau_{A}$ and a number of previous BayeSN analyses only placed upper limits on $\sigma_{R}$ (e.g. Thorp et al., 2021, 2024). In BayeSN, we assume a wide prior $\sigma_{R}\sim\text{Half-}N(0,2^{2})$. This prior is deliberately wide to cover the range of $\sigma_{R}$ values that have been proposed in the literature, and in this work we use BayeSN with its typical priors from previous literature rather than tailoring it to our specific application. However, for a simulated value of $\sigma_{R}=0.5$, with a lack of constraint from the data and such a wide prior an overestimated value of $\sigma_{R}$ is unsurprising. In these cases where $\sigma_{R}$ is overestimated, $N_{95}$ is typically 72 per cent showing that these posteriors do not cover the true simulated value as often as we would expect. With an alternative prior that disfavours large values of $\sigma_{R}$, or more constraint from data we would expect our posteriors to better cover the truth. Indeed, in cases where BayeSN-estimated $\bar{\sigma}_{R}$ values are closer to the true value of 0.5 the values of $N_{95}$ are much closer to 95 per cent.
@@
Table 3: Single $R_{V}$, $E(B-V)$ simulation.
@@
Table 4: Double $R_{V}$, single $E(B-V)$ simulation.
@@
Table 5: Double $R_{V}$, $E(B-V)$ simulation.
@@
Table 6: Double $R_{V}$, $E(B-V)$, and Gaussian $\beta_{\rm SN}$ distribution simulation.
@@ 5.3 Impact of Non-Ia SNe
Table 7 shows the summary of our BayeSN-estimated dust parameters when we simulate with non-Ia contamination. In the case of placing no probability $P_{\rm Ia}$ cuts on our data, we see large biases across all of our parameter estimates, in particular the $\sigma_{Rv}$ components. This is unsurprising given that this sample is dominated by core-collapse supernovae. Indeed, in many cases our BayeSN inference in this case did not converge well given that it was a poor model for this highly contaminated data.
@@
When we impose cuts on $P_{\rm Ia}$, the results improve considerably. Even with a small number of contaminants in the data BayeSN is able to recover true simulated dust parameters with reasonable accuracy. As discussed in Section 2.3, we don’t expect completely unbiased parameter recovery between the two models – even more so with contaminants in the data – but these results are very promising. Differences between the SALT-simulated and BayeSN-inferred dust parameters are overall similar even when including contaminants, after $P_{\rm Ia}$ cuts are applied.
@@
However, it is very noticeable that the BayeSN estimated value for $\tau_{E}$ (HM) does not match the true input value, either with $P_{\rm Ia}$ cuts of 0.5 or 0.9. It is important to note that this is not an issue either with BayeSN or the simulation. In fact, this discrepancy arises because simulated SNe Ia with the highest $E(B-V)$ values are removed by the cut on $P_{\rm Ia}$. Even when using a state-of-the-art photometric classifier such as SuperNNova, significant dust reddening on top of a typical SN Ia template will cause misclassification. This is a very important finding and must be taken into account in future analyses of dust properties on photometrically-classified samples. In practice, this sample with $\mu_{R_{V}}=3$ and $\tau_{E}=0.125$ has a large impact from dust and is more reddened than previous BayeSN analyses have found with real data (e.g. Thorp et al., 2021; Mandel et al., 2022). The number of simulated SNe Ia removed by the photometric classifier is far reduced for the sample with $\mu_{R_{V}}=2$ and $\tau_{E}=0.08$. Nevertheless, our findings demonstrate that great care must be taken to ensure that SNe which are particularly impacted by dust are not misclassified.
@@
The removal of the most highly-reddened SNe means that the sample to which BayeSN was applied had a lower ‘effective’ value of $\tau_{E}$ than the simulated value of 0.125. Across our 50 simulations the mean simulated $E(B-V)$ value was 0.110 and 0.105 for $P_{\rm Ia}$ cuts of 0.5 and 0.9. After the $P_{\rm Ia}$ cuts are applied the $E(B-V)$ distribution will no longer be exactly exponential, but our BayeSN estimated values of $0.108(0.001)$ and $0.102(0.001)$ are close to these ‘effective’ mean $E(B-V)$ values. Considering only the sample of SNe Ia which BayeSN was applied to, it has performed consistently well at recovering the simulated dust parameters.
@@ 6 Application to DES5YR Sample
We test two different BayeSN models on the DES5YR data, for both the Cosmo and BayeDESN samples: one that allows a simple, time-constant peak magnitude difference (intrinsic mag difference) and one that allows for time- and wavelength-dependent changes to the SED as a function of host galaxy mass (intrinsic SED difference).
@@ 6.2 Intrinsic SED difference
$\Delta(g-r)_{\text{peak}}$ $0.042\pm 0.016$ $0.031\pm 0.018$
@@
With regard to host galaxy dust, the choice of an intrinsic mag difference or an intrinsic SED difference does not strongly impact our results. For the Cosmo sample, allowing for an intrinsic SED difference yields $\mu_{R_{V}}$ values of $2.46\pm 0.18$ for the high-mass bin and $3.09\pm 0.32$ for the low-mass bin. For the full BayeDESN sample, these values are $2.53\pm 0.15$ and $3.25\pm 0.22$. The differences between these values are $-0.64\pm 0.35$ and $0.72\pm 0.26$ at significances of $1.8\sigma$ and $2.8\sigma$ respectively. Our inference of $\sigma_{R}$ and $\tau_{A}$ isnot strongly impacted by the choice of parameterisation of the intrinsic mass step.
With either sample, we do not obtain strong constraints on any intrinsic differences around peak. In $g$-band, our inferred differences in intrinsic absolute magnitude at peak are dominated by their uncertainties; for the Cosmo and full BayeDESN samples respectively, we infer $\Delta M_{g,t=0}$ values of $-0.036\pm 0.039$ mag and $-0.019\pm 0.045$ mag. These values are both consistent with 0 but not also inconsistent with an intrinsic magnitude offset of ${\sim}0.04-0.05$ mag, as inferred using our intrinsic mag difference model and in G24.
@@
There is one notable difference between our findings in this work using the DES5YR sample and G24 which analysed DES3YR, Foundation and PanSTARRS Medium Deep. In this work, when considering $g-r$ colour at peak our results suggest that SNe Ia in high-mass galaxies are intrinsically redder than those in low-mass galaxies. For the Cosmo sample, we find $\Delta(g-r)_{\text{peak}}=0.042\pm 0.016$ mag while for the BayeDESN sample the difference is reduced to $\Delta(g-r)_{\text{peak}}=0.031\pm 0.018$. G24 found the opposite, that SNe Ia in high-mass galaxies were intrinsically bluer in $g-r$ colour at peak, with $\Delta(g-r)_{\text{peak}}-0.022\pm 0.010$. As discussed at the beginning of this section, there are fundamental differences between how these samples were observed; DES5YR is a photometric sample, while G24 analysed spectroscopically-confirmed samples. As such, differing conclusions regarding the intrinsic properties of SNe Ia are not surprising; the difference likely arises from greater completeness in the DES5YR sample, although the possible presence of a small number of contaminants in this sample may also partially explain this. Further work is required to make more definitive conclusions about the environmental-dependence of intrinsic colours of SNe Ia.

### v3

@@ 4.2 Impact of non-Ia Contamination
Here, we aim to test the impact of non-Ia contamination on dust inference using BayeSN. To do so, we simulate the Double $R_{V}$, Double E(BV)E(B-V) test with the addition of templates of non-Ia SNe from Vincenzi et al. (2021). We include four classes of non-Ia, following Vincenzi et al. (2024): SN Iax and SN Ia 91bg from Kessler et al. (2019) following revisions from Vincenzi et al. (2019), and core collapse SED templates from Vincenzi et al. (2019) with rates from Strolger et al. (2015) and Shivvers et al. (2017) with the modified Li et al. (2011) luminosity functions from Vincenzi et al. (2021). A full description of the non-Ia templates is available in Vincenzi et al. (2021). We classify our simulated SN light-curves using the SuperNNova classifier from Möller & de Boissière (2020), taking the trained SuperNNova model directly from the DES 5-year analysis, where a further review is available.
@@ 5.2 BayeSN Dust Estimates
The true simulated values input into the Dust2Dust simulations and our BayeSN posterior metrics across all 50 simulacra of our simulations are detailed in Tables 3 through 7.
@@
Firstly, considering $\mu_{R_{V}}$ across all simulated variations the BayeSN estimates are in very close agreement with the truth, within $1\sigma$ in many cases. The most notable discrepancy is $\bar{\mu}_{R}=2.12(0.02)$ for the $\mu_{R_{V}}=2.0$ high-mass bin in the case of the double $R_{V}$, E(BV)E(B-V) simulation with the Gaussian $\beta_{\rm SN}$ distribution, shown in Table 6. The Gaussian $\beta_{\rm SN}$ distribution has no equivalent in BayeSN, and was chosen to test potentially diverging choices in how the two models treat dust; our expectation was that this would be the simulated variant where BayeSN would perform the worst. A discrepancy of ${\sim}0.1$ in $\mu_{R_{V}}$ is small compared to the statistical uncertainty on real inferred $\mu_{R_{V}}$ values; this result could not explain previous disagreements between the findings of BayeSN and Dust2Dust. Additionally, the estimate of $\bar{\mu}_{R}$ in the low-mass bin for this simulation exactly matches the truth of 3.0, showing that this specific model mismatch does not cause particular issues for BayeSN. Our posteriors on $\mu_{R_{V}}$ are relatively well-calibrated considering the inherent model-mismatch in this analysis; $N_{95}$ is close to 95 per cent across all of our simulation variants for $\mu_{R_{V}}$.
@@
While $\sigma_{R}$ is recovered perfectly for our Double $R_{V}$, single E(BV)E(B-V) simulation, it is noticeable that BayeSN estimates of $\sigma_{R}$ are generally further than the true simulated values than for either $\mu_{R_{V}}$ or $\tau_{E}$. This difference is not large and could not explain historic differences between results of SALT-based analyses and BayeSN. For simulations with $\mu_{R_{V}}=3.0$, $\bar{\sigma}_{R}$ is typically overestimated, the most notable example being 0.77(0.04) for simulated $\sigma_{R}=0.5$. For simulations with $\mu_{R_{V}}=2.0$, however, $\bar{\sigma}_{R}$ is typically slightly underestimated, the most notable example being 0.41(0.02) for simulated $\sigma_{R}=0.5$. There are a few reasons why this might be the case. The dust contribution to the $c_{\rm int}$ term in the Dust2Dust simulations (as discussed in Section 2.3) may be impacting the inferred value of $\sigma_{R}$. Alternatively, this might result from the influence of the prior on $\sigma_{R}$ posteriors. $\sigma_{R}$ is typically less constrained by data than either $\mu_{R_{V}}$ or $\tau_{A}$ and a number of previous BayeSN analyses only placed upper limits on $\sigma_{R}$ (e.g. Thorp et al., 2021, 2024). In BayeSN, we assume a wide prior $\sigma_{R}\sim\text{Half-}N(0,2^{2})$. This prior is deliberately wide to cover the range of $\sigma_{R}$ values that have been proposed in the literature, and in this work we use BayeSN with its typical priors from previous literature rather than tailoring it to our specific application. However, for a simulated value of $\sigma_{R}=0.5$, with a lack of constraint from the data and such a wide prior an overestimated value of $\sigma_{R}$ is unsurprising. In these cases where $\sigma_{R}$ is overestimated, $N_{95}$ is typically 72 per cent showing that these posteriors do not cover the true simulated value as often as we would expect. With stronger constraint on $\sigma_{R}$ from data, or perhaps an alternative prior that disfavours large values of $\sigma_{R}$, we would expect our posteriors to better cover the truth. Indeed, in cases where BayeSN-estimated $\bar{\sigma}_{R}$ values are closer to the true value of 0.5 the values of $N_{95}$ are much closer to 95 per cent.
@@
5.2.1 Disentangling Dust Differences from Intrinsic Variation
One of the main applications of BayeSN within G24, as well as in Section 6 of this work, is to probe the cause of the mass step by disentangling intrinsic differences from differences in dust properties. To test this, we consider a simulation including both differences in dust properties and an intrinsic mass step. As discussed in 4.1, in this case we opt to use the ‘intrinsic mag difference’ model rather than the ‘intrinsic SED difference’ model. This is because the simulation does contain a constant achromatic mass step, and using this model directly infers a mass step parameter which we can compare to the truth. Results of this analysis are presented in Table 7.
These results demonstrate that BayeSN can indeed disentangle an intrinsic mass step from differences in dust, even when applied to simulations based on a different model. For a true simulated mass step of 0.05 mag BayeSN infers a step of 0.043 mag on average across all 50 simulacra. While this is slightly offset from the truth, small offsets are not surprising given the difference between the model used for simulation and for inference. It appears that this offset in the mass step is instead captured by a slight offset in the inferred $\tau_{E}$ for SNe Ia in low-mass galaxies. Additionally, the true simulated intrinsic mass step is well covered by the posteriors obtained from BayeSN with $N_{95}=98$ per cent across our 50 simulacra. In fact, the small bias that we observe is comparable to that obtained in a fully SALT-based analysis where the same model was used for both simulation and inference (Popovic et al., 2021).
Overall, BayeSN perfectly recovers the true simulated $\mu_{R_{V}}$ values while finding a significantly non-zero intrinsic mass step between the two populations. This further validates findings presented in G24 and Section 6 of this work that there is an intrinsic component of the mass step alongside a potential contribution from dust.
@@ 5.2.1 Disentangling Dust Differences from Intrinsic Variation
Table 3: Single $R_{V}$, E(BV)E(B-V) simulation.
@@
Table 4: Double $R_{V}$, single E(BV)E(B-V) simulation.
@@
Table 5: Double $R_{V}$, E(BV)E(B-V) simulation.
@@
Table 6: Double $R_{V}$, E(BV)E(B-V), and Gaussian $\beta_{\rm SN}$ distribution simulation.
Parameter Input Value Estimated Value (Std. Err.) $N_{95}$
$\mu_{R_{V}}$ (HM) 3 $3.00(0.03)$ 48(96%)
$\sigma_{R_{V}}$ (HM) 0.5 $0.61(0.02)$ 48(96%)
$\tau_{E}$ (HM) 0.08 $0.081(0.001)$ 48(96%)
$\mu_{R_{V}}$ (LM) 2 $2.00(0.03)$ 47(94%)
$\sigma_{R_{V}}$ (LM) 0.5 $0.44(0.04)$ 49(98%)
$\tau_{E}$ (LM) 0.08 $0.075(0.001)$ 48(96%)
$\Delta M_{\text{int}}$ 0.05 $0.043(0.002)$ 49(98%)
Table 7: Double $R_{V}$ simulation with intrinsic mass step.
@@ 5.3 Impact of Non-Ia SNe
Table 8 shows the summary of our BayeSN-estimated dust parameters when we simulate with non-Ia contamination. These simulations followed the double $R_{V}$, E(BV)E(B-V), and Gaussian $\beta_{\rm SN}$ case as presented in Section 5.2 simulated with additional contaminants. In the case of placing no probability $P_{\rm Ia}$ cuts on our data, we see large biases across all of our parameter estimates, in particular the $\sigma_{Rv}$ components. This is unsurprising given that this sample is dominated by core-collapse supernovae. Indeed, in many cases our BayeSN inference in this case did not converge well given that it was a poor model for this highly contaminated data. For our $P_{Ia}>0$ cut, we find a contamination of 20%; this decreases to 1% for the $P_{Ia}>0.5$ and $P_{Ia}>0.9$ cuts. [^4] The choice of a cut of 0.5 or 0.9 has little impact on the selected sample as few objects have $P_{Ia}$ values between 0.5 and 0.9.
@@
When we impose cuts on $P_{\rm Ia}$ [^5] The ideal solution would be to incorporate $P_{\rm Ia}$ within the likelihood in BayeSN, rather than to cut on $P_{\rm Ia}$. This would require a mixture model treatment of the populations of SNe Ia and contaminants, similar to the Bayesian Estimation Applied to Multiple Species (BEAMS; Kunz et al., 2007, 2013) method. This is planned future work for BayeSN., the results improve considerably. Even with a small number of contaminants in the data BayeSN is able to recover true simulated dust parameters with reasonable accuracy. As discussed in Section 2.3, we don’t expect completely unbiased parameter recovery between the two models – even more so with contaminants in the data – but these results are very promising. Differences between the SALT-simulated and BayeSN-inferred dust parameters are overall similar even when including contaminants, after $P_{\rm Ia}$ cuts are applied.
@@
However, it is very noticeable that the BayeSN estimated value for $\tau_{E}$ (HM) does not match the true input value, either with $P_{\rm Ia}$ cuts of 0.5 or 0.9. It is important to note that this is not an issue either with BayeSN or the simulation. In fact, this discrepancy arises because simulated SNe Ia with the highest E(BV)E(B-V) values are removed by the cut on $P_{\rm Ia}$. Even when using a state-of-the-art photometric classifier such as SuperNNova, significant dust reddening on top of a typical SN Ia template will cause misclassification. This is a very important finding and must be taken into account in future analyses of dust properties on photometrically-classified samples. In practice, this sample with $\mu_{R_{V}}=3$ and $\tau_{E}=0.125$ has a large impact from dust and is more reddened than previous BayeSN analyses have found with real data (e.g. Thorp et al., 2021; Mandel et al., 2022). The number of simulated SNe Ia removed by the photometric classifier is far reduced for the sample with $\mu_{R_{V}}=2$ and $\tau_{E}=0.08$. Nevertheless, our findings demonstrate that great care must be taken to ensure that SNe which are particularly impacted by dust are not misclassified.
@@
The removal of the most highly-reddened SNe means that the sample to which BayeSN was applied had a lower ‘effective’ value of $\tau_{E}$ than the simulated value of 0.125. Across our 50 simulations the mean simulated E(BV)E(B-V) value was 0.110 and 0.105 for $P_{\rm Ia}$ cuts of 0.5 and 0.9. After the $P_{\rm Ia}$ cuts are applied the E(BV)E(B-V) distribution will no longer be exactly exponential, but our BayeSN estimated values of $0.108(0.001)$ and $0.102(0.001)$ are close to these ‘effective’ mean E(BV)E(B-V) values. Considering only the sample of SNe Ia which BayeSN was applied to, it has performed consistently well at recovering the simulated dust parameters.
@@ 6 Application to DES5YR Sample
We test two different BayeSN models on the DES5YR data, for both the Cosmo and BayeDESN samples: one that allows a simple, time-constant peak magnitude difference (intrinsic mag difference) and one that allows for time- and wavelength-dependent changes to the SED as a function of host galaxy mass (intrinsic SED difference). Both of these models jointly infer differences in dust properties alongside intrinsic differences between each environment.
@@ 6.2 Intrinsic SED difference
$\Delta(g-r)_{\text{peak}}$ $042\pm 0.016 0.0$$031\pm 0.018$
@@
With regard to host galaxy dust, the choice of an intrinsic mag difference or an intrinsic SED difference does not strongly impact our results. For the Cosmo sample, allowing for an intrinsic SED difference yields $\mu_{R_{V}}$ values of $2.46\pm 0.18$ for the high-mass bin and $3.09\pm 0.32$ for the low-mass bin. For the full BayeDESN sample, these values are $2.53\pm 0.15$ and $3.25\pm 0.22$. The differences between these values are $0.64\pm 0.35$ and $0.72\pm 0.26$ at significances of $1.8\sigma$ and $2.8\sigma$ respectively. Our inference of $\sigma_{R}$ and $\tau_{A}$ is not strongly impacted by the choice of parameterisation of the intrinsic mass step.
With either sample, we do not obtain strong constraints on any intrinsic differences around peak. In $g$-band, our inferred differences in intrinsic absolute magnitude at peak are dominated by their uncertainties; for the Cosmo and full BayeDESN samples respectively, we infer $\Delta M_{g,t=0}$ values of $0.036\pm 0.039$ mag and $0.019\pm 0.045$ mag. These values are both consistent with 0 but also an intrinsic magnitude offset of ${\sim}0.04-0.05$ mag, as inferred using our intrinsic mag difference model and in G24.
@@
There is one notable difference between our findings in this work using the DES5YR sample and G24 which analysed DES3YR, Foundation and PanSTARRS Medium Deep. In this work, when considering grg-r colour at peak our results suggest that SNe Ia in high-mass galaxies are intrinsically redder than those in low-mass galaxies. For the Cosmo sample, we find $\Delta(g-r)_{\text{peak}}=0.042\pm 0.016$ mag while for the BayeDESN sample the difference is reduced to $\Delta(g-r)_{\text{peak}}=0.031\pm 0.018$. This is inferred simultaneously with the reduced intrinsic peak luminosity difference discussed above, potentially suggesting that requirements for achromatic luminosity offsets can be relaxed by introducing chromatic color offsets, though large uncertainties in peak luminosity offset demonstrate we cannot draw strong conclusions at present. Following, G24 found that SNe Ia in high-mass galaxies were intrinsically bluer in grg-r colour at peak, with $\Delta(g-r)_{\text{peak}}=-0.022\pm 0.010$. As discussed at the beginning of this section, there are fundamental differences between how these samples were observed; DES5YR is a photometric sample, while G24 analysed spectroscopically-confirmed samples. The difference likely arises from greater completeness in the DES5YR sample, although the possible presence of a small number of contaminants in this sample may also partially explain this. Further work is required to make more definitive conclusions about the environmental-dependence of intrinsic colours of SNe Ia.